OpenStudy (anonymous):

explain and solve 2x-5=8

OpenStudy (anonymous):

2x - 5 = 8 2x = 8 + 5 2x=13 x = 2.6

OpenStudy (anonymous):

13 divided by 2 lol my bad 6.5